OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = x2 + x – 6?

OpenStudy (anonymous):

these are the given answers x–intercepts: (3, 0) and (2, 0) y–intercept: (0, 6) x–intercepts: (3, 0) and (–2, 0) y–intercept: (0, –6) x–intercepts: (–3, 0) and (2, 0) y–intercept: (0, –6) x–intercepts: (–3, 0) and (2, 0) y–intercept: (0, 6)

OpenStudy (lalaly):

for the x intercept, set y=0 x^2+x-6=0 (x+3)(x-2)=0 x=-3 and x=2 so the xintercepts are (-3,0) and (2,0) for the y intercept set x=0 y=-6 so the y intercept is (0,-6)

OpenStudy (anonymous):

i have to choose from one of the above lalaly

OpenStudy (lalaly):

the third one,

OpenStudy (anonymous):

oh ok thanx